Coupon = 10% per year, paid Semi-annually matwity = 12 Years * Face value = $1000 required rate of retwas rur: 8% Price :?
Answer #1
Solution:
Bond price =$1,152.47
Working Notes:
Bond Price = Periodic Coupon Payments x Cumulative PVF @ periodic YTM (for t= to t=n) + PVF for t=n @ periodic YTM x Face value of Bond
Coupon Rate = 10%
Annual coupon = Face value of bond x Coupon Rate = 1,000 x 10% = $100
Semi annual coupon = Annual coupon / 2 = $100/2=$50
YTM= 8% p.a (annual)   = required rate of return
Semi annual YTM= 8%/2 = 4%
n= no.of coupon = No. Of years x no. Of coupon in a year
= 12 x 2 = 24
Bond Price = Periodic Coupon Payments x Cumulative PVF @ periodic YTM (for t= to t=n) + PVF for t=n @ periodic YTM x Face value of Bond
= $50 x Cumulative PVF @ 4% for 1 to 24th + PVF @ 4% for 24th period x 1,000
= 50 x 15.24696314 + 1000 x 0.390121474
=$1152.469631
=1152.47
Cumulative PVF @ 4 % for 1 to 24th is calculated = (1 - (1/(1 + 0.04)^24) ) /0.04 = 15.24696314
PVF @ 4% for 24th period is calculated by = 1/(1+i)^n = 1/(1.04)^24 =0.390121474
Please feel free to ask if anything about above solution in comment section of the question.